# An FPGA-based Versatile and Digitalized Method for Pulse Laser Repetition Frequency Locking

### v1

@@ II-D Feedback Loop
In the feedback loop, the amplified frequency difference serves as the error signal for laser stabilization. It is first processed by an LPF to remove high-frequency noise. A PID controller then computes the difference between the error signal and the settled frequency offset. This enables flexible locking of the laser to different relative frequencies. The calculated error is converted to an output voltage by the DAC, and the voltage is amplified by a high-voltage amplifier before being applied to the piezoelectric transducer (PZT) to adjust the laser’s frequency. As all feedback modules are digital, the system can be easily configured to support different feedback devices such as the PZTs and the servos.
@@ III-A Implementation
Figure 4: Block diagram of the error amplification module.
Figure 5: Function modules of FPGA logic firmware. NCO: numerically controlled oscillator.
@@
To evaluate the practicability and the performance of the proposed method, a laser repetition frequency locking system was designed, including the EAM and a high-speed digital FPGA board. Fig.4 shows the block diagram of the EAM and the function modules of the FPGA logic firmware are shown in Fig.5. The laser used in the experiment has a repetition frequency of 50 MHz. Considering that the CFG and CFM used in the system can operate below 2 GHz, a 20x frequency multiplication was applied in the EAM. This effectively amplifies the frequency error by a factor of 20 for subsequent measurements. Fig.6 shows the annotated photograph of the high-speed digital FPGA board. A Xilinx Kintex-7 FPGA is used as the central control unit, with the platform’s digital signal control entirely managed by the programmable digital logic within the FPGA. The power supply section includes both analog and digital circuits to generate the main power supply. Communication between the computer and the FPGA is facilitated through a 1G Ethernet interface. The PLL based CFG provides clock signals to the FPGA, generates ADC sampling signals, and supplies the local oscillator signals for external analog mixing. Additionally, the system includes two CFMs to support multiple channel frequency stabilization. The platform integrates a four-channel high-speed ADC with a resolution of 16 bits and a sampling rate of 100 MS/s. A high-precision DAC, connected via an expansion interface, features a resolution of 20 bits and an output frequency of 1 MS/s. Multiple external expansion interfaces are available for connecting additional DACs and other devices such as DDS. All input and output channels are equipped with $50-\Omega$ Sub-Miniature Version A (SMA) terminations. A 4-GB DDR3 memory provides sufficient data storage capacity and memory bandwidth.
@@
The DMTD frequency extraction method as well as the feedback is deployed in the FPGA for frequency locking. What’s more, a Lock status Detector module can detect the frequency difference in real-time, outputting a locked signal when the frequency error remains within a specified range for a certain period. The Ethernet Interface allows configuration of the EAM and DFLM by computer. Configurable parameters include the multiplication factor, CFG output frequency, frequency measurement period in the Frequency Extraction module, the frequency offset between the laser and reference, low-pass filter parameters, PID parameters, and DAC feedback speed. The EAM is adaptable to different laser repetition frequencies, and the backend feedback module can be configured to accommodate various feedback devices such as PZTs and servos.
@@
This system is suitable for locking a single laser and can be expanded to lock dual laser combs by adding more DACs. The reprogrammable FPGA chip offers excellent flexibility and scalability to meet various requirements. Furthermore, the proposed method offers low latency and high precision, making it suitable for various implementations. By integrating higher-precision ADCs and faster DACs, the system’s performance can be further improved. Additionally, coordinating the use of high-speed and low-speed DACs may enhance locking performance.
@@ III-B DFLM Noise Floor
Fig.7 illustrates the setup for measuring the DFLM noise floor. A reference source is split using a power divider and fed into both channel of the DFLM module through equal-length RF cables. The data is then transmitted via Ethernet Interface to the local PC for stability analysis. Since the two input sine wave signals originate from the same source, their phase difference remains at zero, making this configuration ideal for measuring the system’s noise floor. The reference source used is a rubidium atomic clock, the SAFRAN LPFRS-01, with a nominal frequency of 10 MHz and a stability of $2\times 10^{-11}$ at 1 second. The sampling interval was set to 100 kHz, with a total measurement duration of 400 seconds.
Figure 8: Time deviation in DMTD method and DQD method.
@@
The precision of the DMTD method described in this paper was compared with the DQD method. Both methods use the same external test circuit structure, but the DMTD frequency extraction module is replaced with the DQD frequency extraction module in DQD noise floor test. Fig.8 presents the measured DFLM noise floor, displaying the results both before and after filtering for both methods. The figure shows the unfiltered noise floor and the noise floor after applying a 100 Hz low-pass filter for both methods. The DMTD method demonstrated better noise floor both before and after filtering. After filtering, the DMTD method achieved a noise floor of approximately $1.58\times 10^{-13}@1s$, while the traditional DQD method only reached $9.16\times 10^{-12}@1s$. Additionally, the noise floor of the phase noise analyzer (Microchip model 53100A) are also shown in the figure, taken at a 1 kHz measurement period with a 500 Hz low-pass filter. The analyzer’s noise floor was $7.96\times 10^{-14}$ at a gate time of 1s.
@@
The results demonstrate that the DFLM using the DMTD method achieves a noise floor comparable to that of a commercial phase noise analyzer, offering more precise frequency measurements.
III-C Laser Stability Experiment
@@
Figure 9: Laser repetition frequency offset and Allan deviation before and after locking.
@@
In a standard laboratory environment, a laser frequency locking experiment was performed using a custom-built mode-locked femtosecond fiber laser. Fig.10 shows the scene of the laser frequency locking experiment. The laser operates at a center wavelength of 1550 nm, with a repetition rate of 50 MHz and a tunable range of ±400 Hz, delivering an average optical power of 100 mW. Two 4 mW optical signals were extracted via a fiber coupler and connected to PDs for optical-to-electrical conversion. One signal was directed to the system for frequency locking, while the other was sent to a commercial phase noise analyzer to measure the frequency difference.
@@
The laser repetition frequency was compared to a rubidium atomic clock to generate the error signal. A 500 Hz LPF was applied to the error signal to remove high-frequency noise, and the filtered signal was then sent through a DAC to control the laser repetition frequency by adjusting the PZT. The feedback frequency was set to 1 kHz, which was limited by the PZT.
@@
To evaluate the frequency locking performance and long-term stability, a Microchip phase noise analyzer (model 53100A) was used to measure the frequency difference. Both the reference for frequency locking and the analyzer were connected to a high-stability rubidium atomic clock with a nominal stability of $2\times 10^{-11}@1s$. The gate time was initially set to 0.001 seconds due to the large free-running frequency difference, and later adjusted to 1 second after locking.
@@
Data were collected for two hours both before and after locking. The results are displayed in Fig.9. The left figure shows the laser’s frequency offset before and after locking, demonstrating a significant reduction in frequency fluctuations once locked. On the right, the figure presents the Allan deviation, where the laser’s short-term Allan deviation at 1 second improved from $6.05\times 10^{-8}$ before locking to $1.04\times 10^{-11}$ after. For long-term stability, the laser achieved a stability of $1.12\times 10^{-12}$ at an averaging time of 10 seconds after locking, a substantial improvement from the before-locking stability of $1.51\times 10^{-7}$. Additionally, the right figure shows the noise floor of the phase noise analyzer used in the measurement, as well as results obtained by measuring two different atomic clocks with the same nominal stability. The locked laser exhibited short-term stability comparable to that of the atomic clock, while its long-term stability exceeded that of the atomic clock.
III-D VCO Stability Experiment
Due to the limited stability as well as the feedback bandwidth of our custom-built laser in the Laser Stability experiment, we were unable to fully demonstrate the system’s precision and performance. To address this, the VCO Stability experiment was introduced. In this setup, a VCO was used to simulate the laser’s behavior, as it shares similar characteristics—namely, a stable output frequency and voltage-controlled adjustability. The VCO offers higher feedback bandwidth and stability compared to our laser, making it an effective benchmark for evaluating the system’s performance under current experimental conditions.
The VCO used in this test is the AOCJY1 model from Abracon LLC, with a nominal frequency of 10 MHz, a frequency adjustment range of ±100 Hz, and a nominal stability of $1\times 10^{-9}@1s$. The system was locked using the rubidium atomic clock mentioned earlier as the high stability reference source, which has a nominal stability of $2\times 10^{-11}@1s$. A 300 Hz low-pass filter was employed to filter the error signal, eliminating high-frequency noise, while the feedback frequency was set to 10 kHz.
Figure 11: Fig(a) illustrates the frequency offset of the VCO relative to the reference frequency over 8000 seconds, showing results both before and after locking, as well as the frequency control portion. Fig(b) offers a zoomed-in view of the frequency fluctuations after locking, with a standard deviation measured at 0.0077 mHz. Fig(c) shows the Allan variance before and after locking, with the green line representing the before-locking state and the red line depicting the after-locking state.
The same phase noise analyzer (model 53100A) was used to measure the frequency difference, with the gate time set to 1 second. The rubidium atomic clock was again used as the reference source for the system and the phase noise analyzer. The total measurement time was 6 hours, with 8000 seconds selected for display. As shown in Fig.11(a), from 0 to 836 seconds, the VCO operated in free-run mode. At 837 seconds, the FPGA initiated frequency locking, successfully locking the VCO to the reference frequency within 1 second, maintaining the lock for the remainder of the test. At the end of the figure, the system’s frequency control capability is demonstrated by switching between two different frequency offsets, achieving real-time and precise frequency locking. Unlike traditional phase-locked loops, the system allows for arbitrary frequency locking within the allowable precision range, offering excellent frequency stability with minimal parameter adjustment required.
Fig.11(b) provides a zoomed-in view of the frequency fluctuations after locking, with a standard deviation of 0.0077 mHz. Fig.11(c) presents the Allan deviation before and after locking, with the green line indicating the before-locking state and the red line showing the after-locking state. The short-term Allan deviation at 1 second for the locked VCO frequency reached $9.48\times 10^{-13}$, a significant improvement over the $2.73\times 10^{-10}$ observed before locking. In terms of long-term stability, the VCO exhibited a stability of $1.19\times 10^{-9}$ at an averaging time of 10 seconds before locking, which improved to $9.58\times 10^{-14}$ after locking. The locked VCO demonstrated exceptional stability, confirming the system’s accuracy and reliability.
@@ IV Conclusion
In this paper, we have presented an FPGA-based versatile and digitalized method for pulse laser repetition frequency locking. This method can realize precise frequency measurement, providing reliable frequency lock with minor parameter adjustments using DMTD techniques. The FPGA-based architecture as well as the EAM enables flexibility and adaptability, allowing the system to be reconfigured for various feedback devices and laser frequencies. We design a high-speed FPGA board to evaluate the practicability and the performance of the method.
The results from the Laser Stability experiment demonstrate the system’s capability to significantly reduce frequency fluctuations, achieving a long-term Allan deviation improvement from $1.51\times 10^{-7}$ to $1.12\times 10^{-12}$ at 10 second. The VCO Stability experiment further confirms the system’s precision, with a long-term Allan deviation of $9.58\times 10^{-14}$ at 10 second after locking, making it highly suitable for applications requiring long-term stability.
All these demonstrate that the proposed laser repetition frequency locking method offers a precise, reconfigurable, and compact solution for laser frequency stabilization, with significant potential for future enhancements in both hardware and software.

### v2

@@ II-D Feedback Loop
In the feedback loop, the amplified frequency difference serves as the error signal for laser stabilization. This signal is first processed by an LPF to eliminate high-frequency noise outside the feedback bandwidth. The error signal is then passed to a PID controller, which computes the difference between the error signal and the settled frequency offset, allowing flexible locking of the laser to various frequencies. The calculated error is converted into an output voltage by a DAC, then amplified by a high-voltage amplifier. This voltage is applied to a piezoelectric transducer (PZT) to adjust the laser’s frequency. As all feedback components are digital, the system can be easily configured to support different feedback devices, such as PZTs and servos.
@@ III-A Implementation
Figure 4: Block diagram of the digitized error amplification module.
Figure 5: Function modules of FPGA logic firmware.
@@
The proposed electronic prototype includes an EAM and a high-speed digital FPGA board. Fig.4 shows the block diagram of the EAM. The EAM utilizes the LMK04610 chip from Texas Instruments as the CFG and CFM. This chip can handle input signals from 5 to 500 MHz with output frequencies reaching up to 2 GHz, which is adequate for various laser repetition frequency. BALUNs are employed for signal conversion, minimizing external noise[40]. The CFG provides clock signals to the FPGA, generates ADC sampling signals, and supplies the local oscillator signals for mixing. Based on the sampling rate of the system’s ADC, the EAM generates signals matching the ADC’s sampling frequency.
@@
The function modules of the FPGA logic firmware are shown in Fig.5. The ADC-based-DMTD, frequency extraction method as well as the feedback loop are deployed in the FPGA for frequency locking. What’s more, a lock status detector module can detect the frequency difference in real-time, outputting a locked signal when the frequency error remains within a specified range for a certain period. The Ethernet interface allows configuration of the EAM and DFLM by computer. Configurable parameters include the multiplication factor, CFG output frequency, frequency measurement period in the Frequency Extraction module, the frequency offset between the laser and reference, low-pass filter parameters, PID parameters, and DAC feedback speed. The EAM is adaptable to different laser repetition frequencies, and the backend feedback module can be configured to accommodate various feedback devices such as PZTs and servos.
Fig.6 shows the annotated photograph of the high-speed digital FPGA board. The digital frequency locking board fits within a $20\times 15\times 5$ $cm^{3}$ volume and can be used to lock two independent lasers. A Xilinx Kintex-7 FPGA is used as the central control unit, with the platform’s digital signal control entirely managed by the programmable digital logic within the FPGA. The power supply section includes both analog and digital circuits to generate the main power supply. Communication between the computer and the FPGA is facilitated through a 1G Ethernet interface. The board contains a CFG and two CFMs to support multiple channel frequency stabilization. The platform integrates a four-channel high-speed ADC with a resolution of 16 bits and a sampling rate of 100 MS/s. A high-precision DAC, connected via an expansion interface, features a resolution of 20 bits and an output frequency of 1 MS/s. Multiple external expansion interfaces are available for connecting additional DACs and other devices such as DDS. All input and output channels are equipped with 50 $\Omega$ Sub-miniature version A (SMA) terminations. A 4-GB DDR3 memory provides sufficient data storage capacity and memory bandwidth.
@@
This system is suitable for locking a single laser and can be expanded to lock dual laser combs by adding more DACs. The reprogrammable FPGA chip offers excellent flexibility and scalability to meet various requirements. By integrating higher-precision ADCs and faster DACs, the system’s performance can be further improved. Additionally, coordinating the use of high-speed and low-speed DACs may enhance locking performance.
@@ III-B DFLM Noise Floor
Fig.7 illustrates the setup for measuring the DFLM noise floor. A reference source is split using a power divider and fed into both channel of the DFLM module through equal-length RF cables. The data is then transmitted via Ethernet Interface to the local PC for stability analysis. Since the two input sine wave signals originate from the same source, their phase difference remains at zero[38], making this configuration ideal for measuring the system’s noise floor. The reference source used is a rubidium atomic clock, the SAFRAN LPFRS-01, with a nominal frequency of 10 MHz and a stability of $2\times 10^{-11}$ at 1 second. The sampling interval was set to 100 kHz, with a total measurement duration of 400 seconds.
Figure 8: Time deviation in ADC-based-DMTD method and DQD method.
@@
The precision of the ADC-based-DMTD method described in this paper was compared with the DQD method. Both methods use the same external test circuit structure, but the ADC-based-DMTD frequency extraction module is replaced with the DQD frequency extraction module in DQD noise floor test. Fig.8 presents the measured DFLM noise floor, displaying the results both before and after filtering for both methods. The figure shows the unfiltered noise floor and the noise floor after applying a 100 Hz low-pass filter for both methods. The ADC-based-DMTD method demonstrated better noise floor both before and after filtering. After filtering, the ADC-based-DMTD method achieved a noise floor of approximately $1.58\times 10^{-13}@1s$, while the traditional DQD method only reached $9.16\times 10^{-12}@1s$. Additionally, the noise floor of the phase noise analyzer (Microchip model 53100A) are also shown in the figure, taken at a 1 kHz measurement period with a 500 Hz low-pass filter. The analyzer’s noise floor was $7.96\times 10^{-14}$ at a gate time of 1s.
@@
The results demonstrate that the DFLM using the ADC-based-DMTD method achieves a noise floor comparable to that of a commercial phase noise analyzer, offering more precise frequency measurements.
III-C VCO Stability Experiment
Figure 9: Fig(a) illustrates the frequency offset of the voltage-controlled oscillator (VCO) relative to the reference frequency over 8000 seconds, showing results both before and after locking, as well as the frequency control portion. Fig(b) offers a zoomed-in view of the frequency fluctuations after locking, with a standard deviation (STD) measured at 7.7 µHz root mean square (RMS). Fig(c) shows the Allan variance before and after locking, with the green line representing the before-locking state and the red line depicting the after-locking state. For comparison, the Allan deviation between two independent rubidium atomic clocks is displayed in blue, while the ground noise of the 53100A phase noise analyzer is shown in gray.
To assess the system’s frequency-locking performance, we conducted a stability test using a high-stability voltage-controlled oscillator (VCO). The VCO effectively simulates laser behavior with its stable output frequency and voltage-controlled adjustability. Furthermore, using the VCO minimizes environmental interference, enabling a more controlled evaluation. Its superior short-term stability, low long-term stability, and high feedback bandwidth make it an optimal choice for demonstrating the system’s precision and performance.
The VCO used in this test is the AOCJY1 model from Abracon LLC, with a nominal frequency of 10 MHz, a frequency adjustment range of ±100 Hz, and a nominal stability of $1\times 10^{-9}@1s$. In the EAM, a 100x frequency multiplication was applied, amplifying the VCO frequency to 1 GHz. After down-conversion, the error signal was obtained with an amplification factor of 100. Given the nominal stability of the VCO, a 300 Hz low-pass filter was applied to the error signal in the DFLM to eliminate high-frequency noise, with a feedback frequency set at 100 kHz.
To evaluate the frequency locking performance and long-term stability, a Microchip phase noise analyzer (model 53100A) was used to measure the frequency difference. Both the reference for frequency locking and the analyzer were connected to a high-stability rubidium atomic clock with a nominal stability of $2\times 10^{-11}@1s$. The gate time of the phase noise analyzer is set to 1 second. The total measurement time was 6 hours, with 8000 seconds selected for display. The frequency offset refers to the offset between the measured frequency and the nominal frequency.
As shown in Fig.9(a), from 0 to 836 seconds, the VCO operated in free-run mode. At 837 seconds, the FPGA initiated frequency locking, successfully locking the VCO to the reference frequency within 1 second, maintaining the lock for the remainder of the test. At the end of the figure, the system’s frequency control capability is demonstrated by switching between two different frequency offsets, achieving real-time and precise frequency locking. Unlike traditional PLLs, the system allows for arbitrary frequency locking within the allowable precision range, offering excellent frequency stability with minimal parameter adjustment required. Fig.9(b) provides a zoomed-in view of the frequency fluctuations after locking, with a standard deviation (STD) of 7.7 µHz RMS. Fig.9(c) presents the Allan deviation before and after locking, with the green line indicating the before-locking state and the red line showing the after-locking state. For comparison, the Allan deviation between two independent rubidium atomic clocks is displayed in blue, while the ground noise of the 53100A phase noise analyzer is shown in gray. With locking enabled, the VCO achieved a short-term Allan deviation at 1 second of $9.48\times 10^{-13}$, a significant improvement over the $2.73\times 10^{-10}$ observed before locking. When the averaging time comes to 10 seconds, the VCO exhibited a stability of $1.19\times 10^{-9}$ before locking, which improved to $9.58\times 10^{-14}$ after locking. Even highly stable atomic clocks exhibit inherent random frequency fluctuations. With the VCO locked closely to the reference atomic clock, its frequency precisely followed the reference’s fluctuations, resulting in a measurement stability that surpassed the stability of two independent atomic clocks and approached the analyzer’s baseline noise level.
The locked VCO demonstrated exceptional stability, confirming the system’s accuracy and reliability.
III-D Laser Stability Experiment
@@
Figure 10: (a) Laser repetition frequency offset over 2 hour period. (b) Allan deviation before and after locking.
@@
In a standard laboratory environment, a laser frequency locking experiment was performed using a custom-built mode-locked femtosecond fiber laser. Fig.11 shows the scene of the laser frequency locking experiment. The laser uses a semiconductor saturable absorber mirror (SESAM) and a 9-shaped resonator to generate a femtosecond pulse seed source, achieving passive mode-locking. The repetition rate is tuned by adjusting the cavity length using a PZT. The laser operates at a center wavelength of 1550 nm, with a repetition rate of 50 MHz and a tunable range of ±400 Hz, delivering an average optical power of 100 mW. Two 4 mW optical signals were extracted via a fiber coupler and connected to PDs for optical-to-electrical conversion. One signal was directed to the system for frequency locking, while the other was sent to a commercial phase noise analyzer (model 53100A) to measure the frequency difference.
@@
As the laser used in the experiment has a repetition frequency of 50 MHz, a 20x frequency multiplication was applied in the EAM, which amplifies the frequency error by a factor of 20. In the DFLM, a 500 Hz LPF was applied to the error signal to remove high-frequency noise, and the filtered signal was then sent through a DAC to control the laser repetition frequency by adjusting the PZT. The feedback frequency was set to 1 kHz, which was limited by the PZT.
@@
The system was locked using the rubidium atomic clock mentioned earlier as the high stability reference source, which has a nominal stability of $2\times 10^{-11}@1s$. The same phase noise analyzer (model 53100A) was used to measure the frequency difference and the rubidium atomic clock was also used as the reference source for the phase noise analyzer. The gate time was initially set to 0.001 seconds due to the large free-running frequency difference, and later adjusted to 1 second after locking.
@@
Data were collected for two hours both before and after locking. The results are displayed in Fig.10. The Fig.10(a) shows the laser’s frequency offset before and after locking. The STD of the laser before locking is 45.55 Hz RMS, while the frequency fluctuation stabled at 0.43 mHz RMS after locking, demonstrating a significant reduction in frequency fluctuations once locked. On the right, the Fig.10(b) presents the Allan deviation, where the laser’s short-term Allan deviation at 1 second improved from $6.05\times 10^{-8}$ before locking to $1.04\times 10^{-11}$ after. For the averaging time of 10 seconds, the laser achieved a stability of $1.12\times 10^{-12}$ after locking, a substantial improvement from the before-locking stability of $1.51\times 10^{-7}$. Additionally, the Fig.10(b) shows the noise floor of the phase noise analyzer used in the measurement, as well as results obtained by measuring two different atomic clocks with the same nominal stability.
@@ IV Conclusion
In this paper, we have presented a high-precession, compact electronic prototype for laser repetition frequency locking. The system utilizes a PLL-based CFG and CFM in the EAM to amplify frequency errors. Combined with a high-speed ADC-based DMTD method, it enables precise frequency measurement and reliable frequency locking with minimal parameter adjustments. The FPGA-based architecture as well as the EAM enables flexibility and adaptability, allowing the system to be reconfigured for various feedback devices and laser frequencies.
The results from the VCO Stability experiment confirms the system’s high precision frequency locking capability, with an Allan deviation of $9.58\times 10^{-14}$ at 10 second and an STD of 7.7 µHz RMS after locking, making it highly suitable for applications requiring long-term stability. The Laser Stability experiment demonstrates the system’s capability to significantly reduce frequency fluctuations, achieving an Allan deviation improvement from $1.51\times 10^{-7}$ to $1.12\times 10^{-12}$ at 10 second and an STD of 0.43 mHz RMS after locking.
By incorporating advanced components such as higher-speed, higher-precision ADCs and DACs, or lower-noise PLLs, the system can achieve even greater measurement precision and enhanced feedback bandwidth.
All these demonstrate that the system for laser repetition frequency locking offers a precise, reconfigurable, and compact solution for laser frequency stabilization, with significant potential for future enhancements in both hardware and software.